OpenStudy (anonymous):

6x - 8x less than or equal to 20

OpenStudy (dtan5457):

\[6x-8x \le20\] Combine like values \[-2x \le20\] Divide both sides by -2, and switch the sign (since x is negative)

OpenStudy (anonymous):

\[6x-8x \le 20\] this? Come on, the first thing you can think of :)

OpenStudy (anonymous):

@gennaw do you get it or do you need some morehelp?

OpenStudy (anonymous):

I need more help please

OpenStudy (anonymous):

ok so your inequatlity that you wrote above translated into numeric values would be: \[6x - 8x \le 20\]

OpenStudy (anonymous):

do you get this so far?

OpenStudy (anonymous):

Yes, so far. My teacher hasn't gone over this yet but gives it for homework. Is there another step to the answer

OpenStudy (anonymous):

yes this is just the first step, the second step would be to combine the like terms so the terms that we can combine here are 6x and 8x so do you know how to do this?

OpenStudy (anonymous):

so it would like 6x - 8x can you do that? ^^^

OpenStudy (anonymous):

Sorry no. I am guessing you subtract 6 and 8 to get 2

OpenStudy (anonymous):

yes but positive or negative 2x so 6x - 8x = positive or negative 2x?

OpenStudy (anonymous):

negative?

OpenStudy (anonymous):

thats right! so now the equation simplifies to: \[-2x \le 20\]

OpenStudy (anonymous):

so now you have to isolate x. to do this you would have to divide -2 on both sides of the equation. do you know how to do this?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

wait hold on

OpenStudy (anonymous):

-2x ≤ 20 --- --- -2 -2 the inequality sign changes to the opposite direction when you divide by a negative so it would be x ≥ 20 divided by -2

OpenStudy (anonymous):

do you know hhow to solve from here?

OpenStudy (anonymous):

yes, thank you for helping me

OpenStudy (anonymous):

your welcome! :)